Problem #1

  1. The area between Z = 0 and Z = 2.50

  1. The area between Z = 0 and Z = 2.55

  

  1. The area between Z = 0 and Z = 0.45

  

  1. The area between Z = 0 and Z = -1.33

  1. The area between Z = -1.98 and Z = 2.66
Answer #1

1)

Solution :

Using standard normal table,

a)

P(0 < z < 2.50) = P(z < 2.50) - P(z < 0) = 0.9938 - 0.5 = 0.4938

Area = 0.4938

b)

P(0 < z < 2.55) = P(z < 2.55) - P(z < 0) = 0.9946 - 0.5 = 0.4946

Area = 0.4946

c)

P(-1.33 < z < 0) = P(z < 0) - p(z < -1.33) = 0.5 - 0.0918 = 0.4082

Area = 0.4082

d)

P(-1.98 < z < 2.66) = P(z < 2.66) - P(z < -1.98) = 0.9961 - 0.0239 = 0.9722

Area = 0.9722